OpenStudy (lgbasallote):

i have no idea what this means: \[\mathcal L^{-1} \left \{ \frac{1}{(s^2+a^2)(s^2 + b^2)}\right\}, \quad a^2 \ne b^2; \quad ab \ne 0\]

OpenStudy (anonymous):

a and b are just constant

OpenStudy (lgbasallote):

so those are just to say that this is not undefined?

OpenStudy (anonymous):

they are saying neither of them are 0

OpenStudy (anonymous):

and two are not the equal in value

OpenStudy (lgbasallote):

hmm i'll try doing partial fractions normally..

OpenStudy (anonymous):

typing "Apart" is way to do partial fraction in wolf

OpenStudy (anonymous):

Yes, with partial fraction, you get \[ \frac{1}{\left(a^2+s^2\right) \left(b^2+s^2\right)}=\frac{1}{\left(a^2-b^2\right) \left(b^2+s^2\right)}+\frac{1}{\left(b^2-a^2\right) \left(a^2+s^2\right)} \]

OpenStudy (lgbasallote):

how? :O

OpenStudy (lgbasallote):

well i still need to learn how to get the partial fraction of this manually..

OpenStudy (lgbasallote):

usig gaussian i got coefficient of s^3 A + C = 0 coeff s^2 B+ D = 0 coeff s Ab^2 + Ca^2 = 0 constant Bb^2 + Da^2 = 1

OpenStudy (experimentx):

do the usual stuff ...  since the s^2 get's cancelled ... divide the numerator ... to get the previous value.